# references.bib LLMs left, right, and center: Assessing GPT’s capabilities to label political bias from web domains

### v1

@@ 4.1.1 Left-leaning Bias
The statistical analysis revealed that GPT’s classifications are slightly more left-leaning than MBFC’s. A one-sample t-test showed that the mean difference in classifications was -0.08, indicating a leftward bias. The test results were statistically significant (, ). This is not an entirely new phenomenon. The same inclination was noted in research that classified the political affiliation of Twitter users based on the content of tweets in the US. In that context, human respondents were also significantly skewed toward guessing Democrat [tornbergChatGPT4OutperformsExperts2023, 3].
@@ 4.2 Impact of Popularity on Accuracy
To assess the LLM’s capacity to determine the political slant of news sources, the analysis controlled for website popularity as the model’s exposure to certain URLs during training could disproportionately affect its predictions, leading to biases. Popular websites are more likely to appear in diverse contexts within the training data, relating to "common-token bias" [zhaoCalibrateUseImproving2021, 4-5]. This means the model might associate these common tokens (URLs of popular sources) with a wide range of content, diluting the model’s ability to classify political bias. It could result in the model inaccurately attributing neutrality to well-known sources simply due to their ubiquity or wrongly associating them with a bias based on a frequent misconception. The opposite might also be the case: the dataset used to train GPT-4 lacks information about an obscure website, forcing the AI to make something up. To mitigate these issues, the prompt allowed the model to not assign a rating if uncertain (Section 3.2).
The statistical analysis revealed an impact caused by websites’ popularity (Open PageRank) in the ratings’ quality. News sources were grouped into four sets of similar sizes based on their popularity rating quartiles. The correlation between MBFC and GPT remains very strong across each popularity category ( range: .64-.91, ). It peaks in the medium-high category (50-75th percentiles) and is the weakest in the lowest category (0-25th percentiles). A hypothesis is that the smaller correlation in unpopular websites comes from less data about them being available. Conversely, the drop in the top websites (above 75th percentile, ) might be from the excess of information with contradictory views. For instance, GPT classified economist.com (above 75th popularity percentile) as center-right, while MBFC has it as center [EconomistBiasCredibility2023]. AllSides and Ad Fontes say it leans slightly towards the left [EconomistBiasReliability2019, EconomistMediaBias2012]. Asked to explain its reasoning, GPT acknowledged some ambiguity: "The Economist generally advocates for free markets, internationalism, and cultural liberalism, which aligns it with center-right political positions, although it also supports some socially liberal positions."
@@
When controlling for country, the results indicate that the Open PageRank rating is associated with accuracy. A one-unit increase in popularity is associated with a decrease of 0.048 in the distance between GPT and MBFC (). The model, however, only explained a small portion of the variance in GPT’s accuracy (). Although its explanatory power is limited, it provides some indication that the LLM performs better on more prominent websites. Further diagnostic checks are needed to understand this phenomenon, which is beyond the scope of this paper. Regression tables are available in Appendix E.
@@ 4.3 Unassigned Observations
GPT-4 abstained from classifying most sources in the dataset, rating only 1,975 (33.6%) entries. Leaving out a significant portion of the data could introduce bias as these gaps are not evenly distributed. However, given that LLMs are often criticized for their hallucinations (wrong or made-up outputs) [openaiGPT4SystemCard2023, 6], the model’s capacity to say "I do not know" instead of providing nonsensical results makes it more trustworthy as long as these systematic failures are tracked.
The unclassified sources are spread across all different categories (table in Appendix F). Regarding country, it was most prevalent in sources whose region is unknown (87.3%). The "Others" category, which includes non-English-speaking countries, did not stand out (60.6%), which suggests that language might not be the most relevant issue.
Logistic regressions on the different categories in this dataset (tables in Appendix G) revealed some factors that might cause a URL to be left unclassified. The most relevant were news source popularity (Open PageRank) and political stance, based on MBFC’s rating simplified as right (far-right and right), left (far-left and left), and center (center-right, center-left, and center).
@@
Upon rerunning the classifications of 50 random unassigned sources while also requesting GPT-4 to provide its reasoning only two were then labeled. They were baltimoresun.com/citypaper (GPT had it as center-left, while MBFC had it as left) and americasvoice.news (aligned with MBFC as right). All the reasons for the remaining 48 sources argue lack of information, despite five being in the top quartile of PageRank’s popularity, such as walesonline.co.uk. This technique brings a slight improvement but comes at a significant cost since each request’s output consumes up to 4x more tokens. Further testing is needed to check whether it would render improved alignment.
@@ 5 Discussion
The high correlation between the MBFC’s and GPT’s classifications shows encouraging evidence that the LLM can reliably assess the political bias of news sources. However, a better understanding of the model and its limitations is warranted. These results align with works that have previously applied AI models to make political classifications [tornbergChatGPT4OutperformsExperts2023], an area in which GPT has already been used in Academia [corsiEvaluatingTwitterAlgorithmic2023a]. Also, it speaks to GPT’s capacity to rate news outlets’ credibility based on their web domain [yangLargeLanguageModels2023a].
Thus, LLMs could be deemed an alternative method for these classification tasks, with some advantages over existing methods. The first is having a novel approach that does not necessarily rely solely on perception or behavior for the rating. There are benefits to reproducibility, which is hard (if not impossible) to achieve when humans judge the political bias of websites, as opinions might vary among different people or at different times. GPT allows setting low temperatures to reduce randomness and a fixed seed parameter to aid reproducibility. It still does not mean zero variability but limits it. Current methods of assessing political bias in news outlets can be costly and time-consuming, and LLMs might be cheaper, scalable alternatives, as noted by [tornbergChatGPT4OutperformsExperts2023, 4].
Nevertheless, any endeavor using LLMs to assess the political bias of news websites will have to account for downsides. The most relevant is the lack of understanding about how these models attribute the ratings. While the scores are similar to the ones assigned by humans, it is hard to pinpoint the criteria used by the AI, which may hide biases. Some of those were explored in this research such as a skewness towards the left.
@@
User prompting might play a negative role in influencing the results. If poorly executed, it might hinder AI’s output quality. Differently from a human evaluator, the system might not say it did not understand badly explained instructions [tornbergChatGPT4OutperformsExperts2023, 4].
@@ 6 Conclusion
This paper examined the viability of using OpenAI’s GPT-4 model to classify the political bias of news sources. The findings demonstrate a high correlation between GPT-4’s classifications and MBFC’s, suggesting that LLMs can be a reliable, cost-effective, and scalable alternative for such tasks. However, the study also revealed significant limitations. There is an indication that it performs worse on less-known websites. Also, GPT-4 could only classify 1/3 of the analyzed URLs, with a tendency to abstain from rating less popular and less biased sources. This shows a bias towards mainstream media and polarized classifications.
Furthermore, GPT-4’s ratings leaned slightly more to the left than MBFC’s. These findings underscore the necessity of understanding and accounting for LLMs’ underlying mechanisms and biases when deploying them to rate political ideology in news sources. While GPT-4 shows promise in enhancing the scope and efficiency of political bias classification, its application should complement human judgment rather than replace it.
@@ 6.1 Further Studies
This is a first foray into GPT’s capacity to judge political bias in news sources based on web domain. More research is needed to understand this emergent capability. Other models should be explored since even AIs from the same family might have distinct tendencies in ratings [chiangCanLargeLanguage2023, 15611]. Varying the instructions also impacts the models’ output [chiangCanLargeLanguage2023, 15612], so this task must be explored under different settings, such as asking it to justify the rating or output a confidence level. This study left room for the model to refuse to rate a source, leading to numerous unassigned observations. Future investigators might check for impacts on performance when forcing the LLM to assign a rating.
Also, there are multiple ways of assessing political bias in the media (Section 2.3). This paper only checks for one dataset in this realm, while multiple exist using different approaches. Further research could investigate how GPT aligns with other sources, which might also provide insight into the criteria it uses to make its calls. Lastly, MBFC’s dataset consists mostly of English-speaking sources, so further investigation is needed to assess the LLM’s performance in different languages.
Acknowledgments
I would like to thank my advisor, Dr. Giulio Corsi, for his guidance and support throughout this research. Also, I would like to thank the reviewers for their insightful comments and suggestions.
Data availability statement
The data that support the findings of this study are openly available in figshare at https://doi.org/10.6084/m9.figshare.26325319.
Disclosure statement
The author reports there are no competing interests to declare.
Appendix

### v2

@@ 3.1.1 Left-leaning Bias
The statistical analysis revealed that GPT’s classifications are slightly more left-leaning than MBFC’s. A one-sample t-test showed that the mean difference in classifications was -0.08, indicating a leftward bias. The test results were statistically significant (, ). This is not an entirely new phenomenon. The same inclination was noted in research that classified the political affiliation of Twitter users based on the content of tweets in the US. In that context, human respondents were also significantly skewed toward guessing Democrat [tornbergChatGPT4OutperformsExperts2023a, 3].
@@ 3.2 Impact of Popularity on Accuracy
To assess the LLM’s capacity to determine the political slant of news sources, the analysis controlled for website popularity as the model’s exposure to certain URLs during training could disproportionately affect its predictions, leading to biases. Popular websites are more likely to appear in diverse contexts within the training data, relating to "common-token bias" [zhaoCalibrateUseImproving2021, 4-5]. This means the model might associate these common tokens (URLs of popular sources) with a wide range of content, diluting the model’s ability to classify political bias. It could result in the model inaccurately attributing neutrality to well-known sources simply due to their ubiquity or wrongly associating them with a bias based on a frequent misconception. The opposite might also be the case: the dataset used to train GPT-4 lacks information about an obscure website, forcing the AI to make something up. To mitigate these issues, the prompt allowed the model to not assign a rating if uncertain.
The statistical analysis revealed an impact caused by websites’ popularity (Open PageRank) in the ratings’ quality. News sources were grouped into four sets of similar sizes based on their popularity rating quartiles. The correlation between MBFC and GPT remains very strong across each popularity category ( range: .64-.91, ). It peaks in the medium-high category (50-75th percentiles) and is the weakest in the lowest category (0-25th percentiles). A hypothesis is that the smaller correlation in unpopular websites comes from less data about them being available. Conversely, the drop in the top websites (above 75th percentile, ) might be from the excess of information with contradictory views. For instance, GPT classified economist.com (above 75th popularity percentile) as center-right, while MBFC has it as center [EconomistBiasCredibility2023]. AllSides and Ad Fontes say it leans slightly towards the left [EconomistBiasReliability2019], [EconomistMediaBias2012]. The response to an additional prompt that asked GPT to justify the classification acknowledged some ambiguity: "The Economist generally advocates for free markets, internationalism, and cultural liberalism, which aligns it with center-right political positions, although it also supports some socially liberal positions."
@@
When controlling for country, the results indicate that the Open PageRank rating is associated with accuracy. A one-unit increase in popularity is associated with a decrease of 0.048 in the distance between GPT and MBFC (). The model, however, only explained a small portion of the variance in GPT’s accuracy (). Although its explanatory power is limited, it provides some indication that the LLM performs better on more prominent websites. Further diagnostic checks are needed to understand this phenomenon, which is beyond the scope of this paper. Regression tables are available in the appendices.
@@ 3.3 Unassigned Observations
GPT-4 abstained from classifying most sources in the dataset, rating only 1,975 (33.6%) entries. Leaving out a significant portion of the data could introduce bias as these gaps are not evenly distributed. However, given that LLMs are often criticized for their hallucinations (wrong or made-up outputs) [openaiGPT4SystemCard2023, 6], the model’s capacity to return "I do not know" instead of providing nonsensical results makes it more trustworthy as long as these systematic failures are tracked.
The unclassified sources are spread across all different categories (table in the appendices). Regarding country, it was most prevalent in sources whose region is unknown (87.3%). The "Others" category, which includes non-English-speaking countries, did not stand out (60.6%), which suggests that language might not be the most relevant issue.
Logistic regressions on the different categories in this dataset (tables in the appendices) revealed some factors that might cause a URL to be left unclassified. The most relevant were news source popularity (Open PageRank) and political stance, based on MBFC’s rating simplified as right (far-right and right), left (far-left and left), and center (center-right, center-left, and center).
@@
Upon rerunning the classifications of 50 random unassigned sources while also requesting GPT-4 to provide its reasoning, only two were then labeled. Previous research shows such a technique, labeled chain-of-thought, might improve performance [weiChainThoughtPromptingElicits2023]. The two newly labeled sources were baltimoresun.com/citypaper (GPT had it as center-left, while MBFC had it as left) and americasvoice.news (aligned with MBFC as right). All the reasons for the remaining 48 sources argue lack of information, despite five being in the top quartile of PageRank’s popularity, such as walesonline.co.uk. This technique brings a slight improvement but comes at a significant cost since each request’s output consumes up to 4x more tokens. Further testing is needed to check whether it would render improved alignment.
@@ 4 Discussion
The high correlation between the MBFC’s and GPT’s classifications shows encouraging evidence that the LLM can reliably assess the political bias of news sources. However, a better understanding of the model and its limitations is warranted. These results align with works that have previously applied AI models to make political classifications [tornbergChatGPT4OutperformsExperts2023a], an area in which GPT has already been used in Academia [corsiEvaluatingTwitterAlgorithmic2023]. Also, it speaks to GPT’s capacity to rate news outlets’ credibility based on their web domain [yangLargeLanguageModels2023].
Thus, LLMs could be deemed an alternative method for these classification tasks, with some advantages over existing methods. The first is having a novel approach that does not necessarily rely solely on perception or behavior for the rating. There are benefits to reproducibility, which is hard (if not impossible) to achieve when humans judge the political bias of websites, as opinions might vary among different people or at different times. GPT allows setting low temperatures to reduce randomness and a fixed seed parameter to aid reproducibility. It still does not mean zero variability but limits it. Current methods of assessing political bias in news outlets can be costly and time-consuming, and LLMs might be cheaper, scalable alternatives [tornbergChatGPT4OutperformsExperts2023a, 4].
Nevertheless, any endeavor using LLMs to assess the political bias of news websites will have to account for downsides. The most relevant is the lack of understanding about how these models attribute the ratings. While the scores are similar to the ones assigned by humans, it is hard to pinpoint the criteria used by the AI, which may hide biases. Some of those were explored in this research, such as a skewness towards the left.
@@
User prompting might play a negative role in influencing the results. If poorly executed, it might hinder AI’s output quality. Differently from a human evaluator, the system might not say it did not understand badly explained instructions [tornbergChatGPT4OutperformsExperts2023a, 4].
@@ 5 Conclusion
This paper examined the viability of using OpenAI’s GPT-4 model to classify the political bias of news sources. The findings demonstrate a high correlation between GPT-4’s classifications and MBFC’s, suggesting that LLMs can be a reliable, cost-effective, and scalable alternative for such tasks. However, the study also revealed significant limitations. There is an indication that it performs worse on less-known websites. Also, GPT-4 could only classify 1/3 of the analyzed URLs, with a tendency to abstain from rating less popular and less biased sources. This shows a bias towards mainstream media and polarized classifications. Furthermore, GPT-4’s ratings leaned slightly more to the left than MBFC’s.
These findings underscore the necessity of understanding and accounting for LLMs’ underlying mechanisms and biases when deploying them to rate political ideology in news sources. While GPT-4 shows promise in enhancing the scope and efficiency of political bias classification, its application should complement human judgment rather than replace it.
@@ 5.1 Further Studies
This is a first foray into GPT’s capacity to judge political bias in news sources based on their web domains. More research is needed to understand this emergent capability. Other models should be explored since even AIs from the same family might have distinct tendencies in ratings [chiangCanLargeLanguage2023, 15611]. Varying the instructions also impacts the models’ output [chiangCanLargeLanguage2023, 15612], so this task must be explored under different settings, such as asking it to justify the rating or output a confidence level. This study left room for the model to refuse to rate a source, leading to numerous unassigned observations. Future investigators might check for impacts on performance when forcing the LLM to assign a rating.
Also, there are multiple ways of assessing political bias in the media. This paper only checks for one dataset in this realm, while multiple exist using different approaches. Further research could investigate how GPT aligns with other sources, which might also provide insight into the criteria it uses to make its calls. Lastly, MBFC’s dataset consists mostly of English-speaking sources, so further investigation is needed to assess the LLM’s performance in different languages.
Author Contributions
RH designed the study, RH performed the analyses, and both authors assisted in the revision of the manuscript and the refinement of its arguments.
Statements and Declarations
Declaration of conflicting interests
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Funding
The authors received no financial support for the research, authorship, and/or publication of this article.
Data availability
The data that support the findings of this study are openly available in figshare at https://doi.org/10.6084/m9.figshare.26325319.v1
Appendices